### v2

# Banyan: Improved Representation Learning with Explicit Structure

Abstract
We present Banyan, a model that efficiently learns semantic representations by leveraging explicit hierarchical structure. While transformers excel at scale, they struggle in low-resource settings. Conversely recent structured models have shown promise as efficient learners, but lack performance. Banyan bridges this gap with two key innovations: an entangled hierarchical tree structure and diagonalized message passing, enabling it to outperform larger transformer models with just 14 non-embedding parameters. It excels in low-resource settings, offering a viable alternative for under-represented languages and highlighting its potential for efficient, interpretable NLP in resource-constrained environments.
Machine Learning, ICML
1 Introduction
Semantic representations are foundational for various NLP applications, such as retrieval-augmented generation (RAG) (Lewis et al., 2020), question answering, and summarisation (Abdalla et al., 2023; Wang et al., 2022). They are also crucial for clustering and organising textual data when labelled training data is unavailable. Typically, such representations are generated by large-scale transformer models (Vaswani et al., 2017); highly effective but needing substantial amounts of data and computational resources to train.
An alternative approach draws inspiration from linguistics and cognitive science, incorporating structured compositions—a principle that posits that understanding the semantics of a whole requires knowing the meanings of its parts and the structural rules that determine how they assemble (Chomsky, 1956; Crain & Nakayama, 1987; Pallier et al., 2011; de Marneffe et al., 2006). This principle is highly efficient because novel utterances can be decomposed into familiar components using systematic rules, minimising the need to store individual meanings. It allows humans to learn efficiently from relatively little data and enables effective and efficient learners (Lake et al., 2016; Ito et al., 2022; Wiedemer et al., 2023).
To incorporate such inductive biases into models, the traditional information flow within neural networks needs altering. Instead of relying on implicit processing alone, models must learn representations for atomic components and an explicit computation graph that dictates how these components combine. Additionally, models must learn functions to govern information flow through this graph. Such approaches have demonstrated improved language modelling perplexity at cognitively plausible scales (Hu et al., 2021, 2022), better systematic generalisation (Sartran et al., 2022; Murty et al., 2023), and, especially relevant here, enhanced efficiency in acquiring semantics (Opper et al., 2023).
The Self-StrAE model (Opper et al., 2023) learns representations that explicitly model compositional semantics, and achieves promising performance while requiring minimal resources, both in terms of data and model size. It opened the door to exploring more compute-efficient solutions—particularly valuable for low-resource languages where scaling is often infeasible. However, while innovative, it still falls short compared to large-scale pre-trained transformers, even in languages outside standard pre-training corpora. Here, we introduce Banyan, a model which significantly outperforms Self-StrAE while achieving greater resource efficiency. Our approach involves modifying the structural optimisation process to induce an entangled graph that models global relations between nodes and employs a message passing mechanism using diagonal functions, reducing parameters while enhancing expressiveness.
Banyan, achieves performance comparable to transformer-based baselines and represents a low-cost, viable alternative to transformers for producing representations in low-resource languages, as measured by semantic textual similarity (STS) tasks. By leveraging cognitively inspired inductive biases, our work enables semantic representation learning that rivals or surpasses large-scale pre-trained LLMs—using only 14 non-embedding parameters. This breakthrough offers a new direction for efficient and effective semantic understanding in resource-constrained environments.
2 Background and Related Work
Banyan is a graph neural network, specifically a recursive neural network (RvNN) that learns both structure and representations. We unpack these components below.
Recursive Neural Networks (RvNNs): Like regular recurrent neural networks (RNNs), RvNNs process data by repeatedly applying a function to update their state in sequence. However, instead of relying on temporal ordering (like the sequence of words in a sentence), RvNNs use hierarchical structures, often provided as input—most commonly as a binary tree—and can be applied either bottom-up (from leaves to root) or top-down (root to leaves). They were first introduced by Socher et al. (2011, 2013), inspiring many successor models that vary in how they define the recursive function, including Tree-LSTMs (Tai et al., 2015) and IORNN (Le & Zuidema, 2014).
Learning Structure: RvNNs often require structural input, which limits their flexibility since this structure may not always be available or easily obtainable. To address this, researchers have developed methods to induce structure within the model during recursive computation; using differentiable chart parsing (Drozdov et al., 2019, 2020; Hu et al., 2021, 2022), beam search (Ray Chowdhury & Caragea, 2023), continuous relaxation (Chowdhury & Caragea, 2021; Soulos et al., 2024), and reinforcement learning (Havrylov et al., 2019). However, these methods can struggle with memory issues and require careful tuning of hyperparameters. Here, we adopt a method from Opper et al. (2023) that uses representation similarity to determine how nodes should be merged during computation, which is both computationally efficient and surprisingly effective.
Semantic Representations of Text: Systems like Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) use the distributional hypothesis (Harris, 1954) to model word semantics, which posits that words are defined by the context in which they appear. To learn representations, these models use a fixed context window and predict a missing word in a sequence. While initially effective, this approach has limitations because words can have multiple meanings depending on context. Early encoder-only transformer models produced poor representations (Reimers & Gurevych, 2019), but recent developments with contrastive finetuning (Gao et al., 2021) have become the method of choice for producing semantic representations.
Semantic Representation Learning through Structure: Transformer embeddings have become more successful than static word embeddings due to their ability to handle varying contextual influences. Unlike attention mechanisms in transformers, which route information based on token relationships, some approaches use explicit graphs or structures—such as dependency (Levy & Goldberg, 2014; Vashishth et al., 2019) or constituency (Opper et al., 2023) parses—to determine the focus of context windows. Opper et al. (2023) uses constituency parsers to learn sentence-level embeddings alongside word embeddings, introducing two models: StrAE, which relies on being given its structure, and Self-StrAE, which learns its own structure using representations. The latter model, Self-StrAE, serves as the foundation for Banyan and is described next.
3 Preliminary: Self-StrAE
[…]
Homer
[…]
ate
$\bar{e}_{4}$ $\Omega_{\Psi}$ $w_{3}$
doughnuts
[…]
Homer
[…]
ate
$\underline{e}_{4}$ $\Lambda_{\Gamma}$ $\hat{w}_{3}$
doughnuts
root embedding shared
Figure 1: Self-StrAE operation. Red lines indicate cosine similarity. Shared colours imply shared parameters.
Self-StrAE involves three main components that act over a sequence of tokens $\mathbf{w}=\langle w_{n}\rangle_{n=1}^{N}$: (a) an algorithm for merging tokens based on their similarity, (b) functions for composition and decomposition of embeddings, and (c) an objective that leverages both the induced structure and embeddings. While full details are available in Opper et al. (2023), we provide a brief overview to establish context for our model development (§ 4).
At a high level, Self-StrAE learns representations that define their own structure while being shaped by it. Starting with an initial embedding matrix $\Omega_{\Psi}$, tokens are merged into single embeddings using a composition function $C_{\Phi}$ based on best cosine similarity (e.g., see Figure 1). This process reduces the sequence to a single root embedding while capturing semantic relationships. The resulting merge history forms a binary tree structure, over which the model then operates in reverse by decomposing embeddings at each node using a decomposition function $D_{\Theta}$, to reconstruct the leaf embeddings. Optionally, it can further predict tokens ($\widehat{w}_{n}$) using a dembedding function $\Lambda_{\Gamma}$. Figure 1 illustrates the autoencoding process. During training, tokens that are frequently merged together develop correlated representations, leading the model to learn meaningful compositional semantics. This results in embeddings that reflect both their own structure and the semantic patterns they encode.
More formally, one denotes tokens as the vertices $w_{i}\in\Delta^{V}$ in a $V$-simplex for vocabulary size $V$, and note that the models generates two sets of embeddings—one going up ($\bar{e}$: leaves $\to$ root) and one coming down ($\underline{e}$: root $\to$ leaves). The embeddings are viewed as $e\in\mathbb{R}^{U\times{}K}$ allowing the composition and decomposition functions to act independently over $K$ channels, and be defined as
[…]
To learn this model from data, Opper et al. (2023) derive two objectives. The first is straightforward cross-entropy over reconstructed tokens, for sentence $\mathbf{w}$ and prediction $\widehat{\mathbf{w}}$ as $\mathcal{L}_{\text{CE}}(\mathbf{w},\widehat{\mathbf{w}})=-\frac{1}{N}\sum_{n=1}^{N}w_{n}\cdot\log\widehat{w}_{n}.$ This however, places little constraint on the intermediate nodes in the hierarchical model. To address this, an alternate structural contrastive objective is formulated over a batch of sentences. As up and down trees are structurally identical (modulo edge reversal), it draws together an embedding and its dual on the other tree, while pushing away all other embeddings across the batch, using cosine similarity. Denoting pairwise similarity matrix $A\in\mathbb{R}^{M\times M}$ between up and down embeddings over $M$ nodes in the batch, the objective is: $\mathcal{L}_{\text{CO}}(\bar{e},\underline{e})=\frac{-1}{2M}\sum_{i=1}^{M}\log\left(\sigma_{\!\tau}(A_{i\bm{\cdot}})\,\sigma_{\!\tau}(A_{\bm{\cdot}i})\right)$ with $\sigma_{\!\tau}(\cdot)$ the tempered softmax over the unspecified ($\bm{\cdot}$) dimension.
4 Banyan
[…]
Entangling
[…]
Figure 2: Entangled trees: Example of disjoint trees being transformed into an entangled tree. Internal functions ($C_{\Phi},D_{\Theta},\ldots$) are elided to avoid clutter.
4.1 From trees to entangled trees
Algorithm 1 Banyan: Entangled Compose
0:  Global frontier $\langle(s_{n},e_{n})\rangle_{n=1}^{N}$, compose ($\circ$), concat ()($\diamond$)( ⋄ ), similarity cSim$(e,e^{\prime})$
1:  $\mathcal{A}\leftarrow\langle(s_{n},e_{n})\rangle_{n=1}^{N}$                       $\triangleright$ initialise frontier
2:  $(\mathcal{V},\mathcal{E})\leftarrow(\varnothing,\varnothing)$                          $\triangleright$ initialise graph
3:  while $\exists i:s_{i}\diamond s_{i+1}\not\in_{s}\mathcal{V}$ do
4:     ${i^{\star}}\leftarrow\operatorname*{arg\,max}_{i}\textsc{cSim}(e_{i},e_{i+1})$        $\triangleright$ locate closest pair
5:     $e_{p}=\circ(e_{i^{\star}},e_{i^{\star}+1})$                      $\triangleright$ compose
6:     $\mathcal{V}\leftarrow\mathcal{V}\cup\{(s_{i^{\star}}\diamond s_{i^{\star}+1},e_{p})\}$
7:     $\mathcal{E}\leftarrow\mathcal{E}\cup\{{p}\sim{i^{\star}},{p}\sim(i^{\star}+1)\}$
8:     $\mathcal{J}\leftarrow\{j:(s_{j},s_{j+1})=(s_{i^{\star}},s_{i^{\star}+1})\}$                       $\triangleright$ locate all occurrences of this pair
9:     $\mathcal{A}\leftarrow\mathcal{A}\setminus\{\forall_{j\in\mathcal{J}}\;\mathcal{A}_{j},\mathcal{A}_{j+1}\}$                      $\triangleright$ delete occurrences from those locations
10:     $\mathcal{A}\leftarrow\mathcal{A}\cup_{\mathcal{J}}\{(s_{i^{\star}}\diamond s_{i^{\star}+1},e_{p})\}$                       $\triangleright$ insert composition into those locations return: Graph $(\mathcal{V},\mathcal{E})$
We wish to have composition embeddings amortise over all possible contexts, and simultaneously, all decompositions embeddings to resolve to the same thing. The representation of an entity Lisa should encapsulate everything she could possibly eat. Simultaneously, the average of everything she could eat we should get back to Lisa. Self-StrAE does not explicitly model this behaviour in its structure. Decomposition embeddings of the same entity only interact when we calculate the loss. On top of this, because the loss is taken over the batch, they are actually treated as false negatives to each other. Even though they are terms that ought not be pushed away, the objective ask them to be.
Our innovation is to address both these issues by formulating the process in terms of entangled trees—where entangling describes the transformation of disjoint tree structures into a conjoined graph structure. An example is shown in Figure 2. Here, all instances of “night” and “some are born to” are captured by a single node representing that constituent. We call our model Banyan on account of this entangling, because, like the tree, it can have many roots—consisting of nodes frequently reused across contexts.
Entangling: Constructing an entangled tree given a set of disjoint trees is a relatively straightforward process and is formally specified in Algorithm 1. In contrast to the agglomerative clustering employed in Self-StrAE, here we employ a global frontier spanning all leaf nodes across the given data. The key differences to the prior methods are mainly to do with constructing a graph jointly with progressing the frontier and ensuring that new nodes are never duplicated, for which we employ a node identity $s_{n}$ in addition to the node embedding $e_{n}$.
Incorporating context: Following the entangling of trees described, the model proceeds in a similar vein to Self-StrAE, by composing upwards from leaves to roots (multiple roots corresponding to multiple trees), and then decomposing downwards back to the leaves. With entangled trees, while traversing upwards each node is always composed from the same two children, but on the way back down, things are different as each separate context for a given node provides a different downward embedding. This is shown in Figure 3 focussing on a subgraph of the entangled tree from Figure 2(right). Note that the node in question (in blue) corresponds to the span “some are born to”, and has downward embeddings that incorporate context both from “endless night” and “sweet delight”. This is exactly as desired, as Banyan allows explicit aggregation to derive the downward embedding that resolves over the contexts. For any upward embedding $\bar{e}$ whose span occurs in different contexts $y\in\mathcal{Y}$, the corresponding downward embedding is derived by simply averaging over the different contextual down embeddings; i.e., $\underline{e}=1/\lvert\mathcal{Y}\rvert\sum_{y}\underline{e}^{y}$.
[…]
Figure 3: Upward and downward traversals for a section of the entangled tree from Figure 2.
Effectiveness and efficiency: Beyond the ability to explicitly incorporate context across data, entangled trees also help the contrastive objective by avoiding false negatives since they do not admit duplicate nodes by construction. Furthermore, the lack of duplicate nodes also drastically impacts the memory footprint of the model as one deals with the set of all nodes rather than counting each instance as its own node. These effects becomes more pronounced when entangling a larger set of instances as the likelihood of false negative and duplicates goes up together.
Practical estimation: Given the advantages of entangled trees, one would ideally want to construct it over all the available data—not practically feasible with the exponential growth in dataset sizes. To address this, we construct our model to estimate the given objective by taking steps over batches of data that are of a more manageable size, noting that this estimator is unbiased. To see this is the case, note that entangled trees only affects the downward embeddings directly, and that batching simply means that the resolved embedding is an average over samples instead of over all the data (population)—the sample mean is always an unbiased estimator of the population mean.
4.2 Simplified Message Passing
Complementary to the development of entangled trees to incorporate context, we also explore avenues to improve the message passing with the composition ($C$) and decomposition ($D$) functions. The original formulations Equations 1 and 2 concatenate or split using simple single-layer linear neural networks. These were found to lead to better representations than e.g., Tree-LSTM cells, because they forced the model to conform to the compression order of the structure.
But if all that was required for success is to respect the compression order, then one could possibly do better with a simpler solution that exploits diagonalised functions (Ba et al., 2016)—a crucial component in the resurgence of recurrent neural networks (Peng et al., 2023; Orvieto et al., 2023; De et al., 2024) introducing decayed memory across time. Thus, rather than using linear layers, we now define:
$C(\bar{e}_{i},\bar{e}_{i+1})$ $\!=\!(\bar{e}_{i}\!\cdot\!\sigma(\Phi_{l})+\bar{e}_{i+1}\!\cdot\!\sigma(\Phi_{r}))+\phi$ (3)
$D(\underline{e}_{i})=\bigl{(}$ $\underline{e}_{i}\!\cdot\!\sigma(\Theta_{l})+\theta_{l},\;\;\underline{e}_{i}\!\cdot\!\sigma(\Theta_{r})+\theta_{r}\bigr{)}$ (4)
$\Phi_{l},\Phi_{r},\phi,\Theta_{l},\Theta_{r},\theta_{l},\theta_{r}\in\mathbb{R}^{U}$
with sigmoid non-linearity ($\sigma$) applied to parameters both for numerical stability and to enforce a decayed memory over structure depth. Repeated application of the diagonal composition function will decay the influence of nodes further down in the tree, thereby respecting the compression order of the structure. Also, during composition parent representations can increase in magnitude as they are the sum of the two children. During decomposition child representations will, by necessity, reduce back down in magnitude towards the leaves. In this way the functions further mimic the information flow specified by the entangled trees.
These relatively simple changes have a pretty drastic effect, both in terms of performance (see experiments) as well as memory footprint, with parameters now reduced by a factor of $U$ compared to the functions from Equations 1 and 2
Table 1: Sentence level results for models pretrained on English. Higher is better. Results averaged across four random initialisations. Bolded results indicate no standard deviation overlap. Spearman’s $\rho$ is * 100 following convention.
Model STS-12 STS-13 STS-14 STS-15 STS-16 STS-B SICK-R SemRel Score
Self-StrAE 31.98 $\pm$ 0.58 53.88 $\pm$ 0.68 37.73 $\pm$ 0.70 55.23 $\pm$ 0.58 $5$$\pm$ 0.47 39.53 $\pm$ 1.61 51.78 $\pm$ 0.29 50.05 $\pm$ 0.92 46.59 $\pm$ 0.43
GloVe 31.61 $\pm$ 0.31 21.69 $\pm$ 0.12 27.37 $\pm$ 0.10 40.42 $\pm$ 0.09 29.27 $\pm$ 0.12 28.25 $\pm$ 0.08 50.20 $\pm$ 0.25 41.20 $\pm$ 0.43 33.75 $\pm$ 0.04
+ stopword rm 39.00 $\pm$ 0.57 41.61 $\pm$ 0.19 39.31 $\pm$ 0.18 51.06 $\pm$ 0.35 45.14 $\pm$ 0.14 48.40 $\pm$ 0.07 52.80 $\pm$ 0.04 42.37 $\pm$ 0.13 44.96 $\pm$ 0.10
RoBERTa 42.77 $\pm$ 1.27 51.70 $\pm$ 1.30 45.67 $\pm$ 1.42 63.97 $\pm$ 0.81 59.60 $\pm$ 0.61 39.97 $\pm$ 0.95 52.93 $\pm$ 0.23 52.73 $\pm$ 0.58 51.08 $\pm$ 0.61
+ SimCSE 50.63 $\pm$ 1.45 62.23 $\pm$ 2.51 54.17 $\pm$ 2.10 68.77 $\pm$ 3.00 66.67 $\pm$ 1.40 53.53 $\pm$ 1.18 56.87 $\pm$ 1.16 59.27 $\pm$ 0.93 59.02 $\pm$ 1.45
Banyan 51.20 $\pm$ 0.007 69.10 $\pm$ 0.002 63.20 $\pm$ 0.004 73.20 $\pm$ 0.002 66.60 $\pm$ 0.002 61.50 $\pm$ 0.002 55.50 $\pm$ 0.003 61.60 $\pm$ 0.002 62.70 $\pm$ 0.001
Table 2: Word level results analogous to Table 1.
Model Simlex Wordsim-S Wordsim-R Score
Self-StrAE 13.80 $\pm$ 0.41 54.38 $\pm$ 0.78 52.85 $\pm$ 1.27 40.34 $\pm$ 0.66
GloVe 27.47 $\pm$ 0.25 62.53 $\pm$ 0.42 51.00 $\pm$ 0.56 47.00 $\pm$ 0.38
RoBERTa 29.23 $\pm$ 0.64 61.97 $\pm$ 2.38 46.00 $\pm$ 2.13 45.73 $\pm$ 1.71
Banyan 16.57 $\pm$ 0.02 63.25 $\pm$ 0.03 69.00 $\pm$ 0.01 49.61 $\pm$ 0.02
5 Experiments:
5.1 Warmup: English language evaluation
Goal: We wish to test whether Banyan can efficiently learn semantics. We start by evaluating on English, as there are far more test sets available than for low resource languages.
Evaluation: We want to evaluate how well Banyan learns effective semantic representations. Ideally we want to probe this at different levels of hierarchy, covering both the lexical and sentential level. Our evaluation is unsupervised, both to directly probe the effect of pretraining with the inductive bias, and because this setting has greater parity to what may be expected in a low-resource domain, with few labelled datasets. For these reasons, we turn to a series of tasks which measure correlation between cosine similarity of embedding pairs for two examples and human judgements of their semantic correspondence. On the word level, we use Simlex-999 (Hill et al., 2015) and WordSim-S/R (Agirre et al., 2009). All tasks measure semantics, but do so on differing axes. Simlex measures similarity at the exclusion of relatedness. Wordsim S measures similarity without penalising relatedness. And Wordsim R measures relatedness. On the sentence level, we use STS-12 through 16 (Agirre et al., 2012, 2013, 2014, 2015, 2016), the STS-B (Cer et al., 2017), SICK-R (Marelli et al., 2014) and SemRel (Ousidhoum et al., 2024). In addition to semantic similarity we look at retrieval and classification tasks to ensure embedding robustness. We use the Quora and Arguana retrieval datasets from the BEIR suite (Thakur et al., 2021) alongside MRPC and SST-2 from the GLUE benchmark. For the classification tasks we train a GeLU MLP with a 512D hidden size, and leave models frozen to emphasise the effect of pre-training. A good embedding model should do well across the board.
Baselines: We compare against the Self-StrAE, GloVe embeddings (Pennington et al., 2014) and a RoBERTa (Liu et al., 2019) in the medium configuration from (Turc et al., 2019). Self-StrAE stands as the closest point of comparison to Banyan. Self-StrAE indicates the performance level of structured representation learning lies, as well as any improvements we are able to achieve. GloVe lets us compare to traditional static embeddings. This comparison probes whether our model is learning anything more than just simple bag of word features. To obtain sentence embeddings, we report results using both the simple average of the word embeddings and the average with stopwords removed following (Reimers & Gurevych, 2019). The latter is generally stronger. For RoBERTa, we report results using both the standard model, and again after enhancing RoBERTa through an extra round of contrastive SimCSE training (Gao et al., 2021), as a further STS baseline. Our pooling strategy is mean. To produce static embeddings from RoBERTa to use in lexical evaluation, we follow Bommasani et al. (2020) and average the contextualised representations of all occurrences of the word in the training set. The RoBERTa is intended as a stronger baseline. It has significantly more parameters than Banyan and is able to model meaning in context unlike GloVe.
Hyperparameters and Pre-training Details: For all models we set the embedding size to 256. For Self-StrAE we use the configuration of (Opper et al., 2023) and set embeddings as square matrices (i.e., $K$=16 and $U$=16). For Banyan we set these values to $K$=128 and $U$=2, because the more independent channels we allowed the better the model seemed to perform. We refer the reader to the Appendix for ablations. We also note that because we can perform this reduction in channel size, the number of non-embedding parameters for Banyan drops to just 14, as these are directly proportional to $U$. We trained Self-StrAE and Banyan for 15 epochs (circa 15k steps and sufficient for convergence) using the Adam optimizer (Kingma & Ba, 2015), with a learning rate of 1e-3 for Banyan and 1e-4 for Self-StrAE  using a batch size of 512. To process the graphs we used DGL (Wang et al., 2020). The GloVe baseline was trained for 15 epochs with a learning rate of 1e-3, and a window size of 10. We used the official C++ implementation. RoBERTa medium was trained for 200,000 steps, (10% of which were used for warmup). We used a learning rate of 5e-5, and a linear schedule. Positional embeddings are relative key-query. The configuration for RoBERTa medium is 8 layers, 8 attention heads and 2048 dimensional feedforward layers. We used the Transformers library to implement and train the model (Wolf et al., 2020). For SimCSE training, we used the default parameters and the official implementation for unsupervised RoBERTa training from Gao et al. (2021). As our pre-training corpus we selected the WikiText-103 benchmark dataset (Merity et al., 2016). The RoBERTa and GloVe baselines are trained on the full corpus (103 million tokens), representing the upper-middle end of the level of data scale that might be available for a language. Whereas we trained Self-StrAE and Banyan on a uniform subsample of 10 million tokens, representing the lower end of how many tokens might be available, because these explicit structure models are supposed to be efficient learners.
Results: Results are shown in Tables 1, 2 and 3. On both the word level and sentence level Banyan does much better than Self-StrAE. We ablate the reasons for this in more detail later in the manuscript. Both models suffer on SimLex because they need to model both similarity and relatedness as the latter dictates merge (related concepts often compose together). However, the important thing to note is that the structured models effectively transfer the same performance from the word level to the sentence level. They can take advantage of composition, and transfer the meaning of the parts to understanding the meaning of the whole. The GloVe baseline is good on the word level, but does not generalise to the sentence level as well as the transformer, even when we give it stopword removal. It cannot transfer semantic knowledge seamlessly to different levels of complexity. Banyan can, and is able to achieve comparable or better performance than the SimCSE RoBERTa despite being much smaller and exposed to 10x less pre-training data. This means we have a structured model that remains efficient and cheap, and also effective at representation learning.
Table 3: Sentence level results on retrieval and classification tasks for models pretrained on English.
Quora Arguana SST-2 MRPC
Model NDCG@1 NDCG@10 R@1 R@10 NDCG@1 NDCG@10 R@1 R@10 Acc F1
Self-StrAE 32.88 $\pm$ 0.28 40.02 $\pm$ 4.94 29.59 $\pm$ 0.23 44.77 $\pm$ 0.28 10.96 $\pm$ 0.11 15.48 $\pm$ 0.13 09.96 $\pm$ 0.11 21.48 $\pm$ 0.22 74.67 $\pm$ 0.52 80.34 $\pm$ 0.42
GloVe 29.99 $\pm$ 0.14 35.71 $\pm$ 0.15 26.08 $\pm$ 0.15 43.17 $\pm$ 0.25 06.19 $\pm$ 0.19 12.77 $\pm$ 0.24 06.18 $\pm$ 0.19 24.68 $\pm$ 7.40 75.83 $\pm$ 0.62 81 $\pm$ 0
+ stopword rm 44.41 $\pm$ 0.13 52.54 $\pm$ 0.17 38.78 $\pm$ 0.15 62.15 $\pm$ 0.25 09.89 $\pm$ 0.19 20.27 $\pm$ 0.09 09.89 $\pm$ 0.19 33.00 $\pm$ 0.26 76.50 $\pm$ 1.08 81 $\pm$ 0
RoBERTa 43.26 $\pm$ 0.76 49.97 $\pm$ 0.72 37.67 $\pm$ 0.68 58.78 $\pm$ 0.78 08.18 $\pm$ 0.43 17.60 $\pm$ 0.36 08.18 $\pm$ 0.43 28.85 $\pm$ 0.94 75.68 $\pm$ 0.96 81 $\pm$ 0
+ SimCSE 51.79 $\pm$ 2.12 59.30 $\pm$ 2.10 45.09 $\pm$ 1.60 68.74 $\pm$ 2.01 10.06 $\pm$ 1.27 21.84 $\pm$ 2.23 10.06 $\pm$ 1.27 37.36 $\pm$ 2.16 75.97 $\pm$ 1.08 80.83 $\pm$ 0.24
Banyan 57.83 $\pm$ 0.04 65.78 $\pm$ 0.05 50.19 $\pm$ 0.08 75.80 $\pm$ 0.18 13.21 $\pm$ 0.25 29.28 $\pm$ 0.11 27.41 $\pm$ 0.68 49.60 $\pm$ 0.52 77.2 $\pm$ 0.16 79.51 $\pm$ 0.27
[…]
Model Indonesian Arabic Telugu Marathi Mor. Arabic Kinyarwanda Hausa Afrikaans Spanish Amharic Hindi Score
XLM-R 46.7 31.6 46.3 55.7 17.4 13.2 4.1 56.2 68.9 57.3 52.7 40.92
Llama-3.1 (8B) 53.4 31.1 65.6 63.4 19.4 19.7 6.1 65.4 66.7 64.1 61.7 46.96
Mistral Nemo 50.7 20.1 57 52.3 15.1 16.3 1.8 58.3 66.2 53.2 55.8 40.62
[…]
XLM-R (FT) 47.9 33.6 68.8 75.1 21.6 19.4 14.6 72.6 72.8 59.6 57.6 49.41
Banyan 44.17 43.20 71.13 67.67 52.00 46.1 43.7 78.68 60.95 66.18 61.83 57.78
$\pm$ 1.11 $\pm$ 1.82 $\pm$ 0.91 $\pm$ 0.64 $\pm$ 2.25 $\pm$ 0.32 $\pm$ 1.21 $\pm$ 0.30 $\pm$ 0.76 $\pm$ 0.46 $\pm$ 0.6
5.2 Multilingual Evaluation:
Goal: From the results in English we know that Banyan is an efficient learner: it can produce good representations without requiring large-scale data or compute. This implies potential use for under represented communities, whose languages are not well covered by current NLP approaches. Now we have to test that.
Evaluation: Learning semantic representations for low-resource languages remains an ongoing challenge in NLP. A core problem is not just the lack of training data, but also the lack of evaluation datasets. Recent work by Ousidhoum et al. (2024) has sought to address this issue, providing semantic relatedness test sets for several low resource Asian and African languages. These test sets are evaluated the same as before, comparing the cosine similarity between model embeddings for two sequences with human judgements of their semantic match. As before, evaluation is zero-shot unsupervised, allowing us to evaluate Banyan on Indonesian, Arabic, Telugu, Marathi, Moroccan Arabic, Kinyarwanda, Hausa, Afrikaans, Spanish, Amharic and Hindi—covering a broad spectrum of resource extent. For example, Spanish and Hindi are reasonably well represented, while Moroccan Arabic and Kinyarwanda have extremely little training data.
Baselines: We select XLM-R (Conneau et al., 2019): a transformer encoder trained on 2TB of multilingual data. Llama 3.1 8B (Dubey et al., 2024): a decoder only LLM trained on 15 trillion tokens. Mistral Nemo 12B: a decoder only LLM designed with multi-lingual capacities in mind. In addition we also compare against two specialised embedding models from the sentence transformers range (Reimers & Gurevych, 2019): Mini-LM-L12-V2 and Paraphrase-XLM-R-Multilingual-V1. These are pre-trained transformer encoders that have been finetuned on supervised datasets designed to produce high quality semantic representations. The baselines we select here are emblematic of the kind of models one might reach for in order to embed a corpus. For all models we use mean pooling to produce the sentence representation following (Reimers & Gurevych, 2019; Li & Li, 2024). Finally, for parity we include an XLM-R baseline which is finetuned on the same corpora.
Banyan Pre-training and Hyperparameters: For Afrikaans, Spanish and Amharic we obtained corpora from Leipzig Corpora Collection [^1] For Spanish and Hindi we select the mixed corpus and uniformly subsample to reduce size to $\approx$10M tokens. (Goldhahn et al., 2012). For Amharic we utilised a MiT licenced pre-training set of 1 million sequences available on the Huggingface hub at this link. Kinyarwanda and Hausa data was sourced from Opus (Nygaard & Tiedemann, 2003). Each dataset consists of roughly 10 million tokens. We utilise a pre-trained BPE tokenizer for each language from the BPEMB Python package (Heinzerling & Strube, 2018). For the model hyperparameters we keep all the settings from the experiments on English. For XLM-R we finetune for up to 100k steps with early stopping, using a linearly scheduled learning rate of 5e-5 with 10% of steps as warmup. XLM-R runs at batch size 128 across 4xA40 45gb cards.
Results: See Table 4. In Spanish, a well-resourced language with high coverage, the transformer baselines almost all outperform Banyan. However, as languages become lower resourced the picture changes, and Banyan outperforms or is comparable to the baselines. This even includes the multilingual XLM-R that has undergone supervised training. While finetuning XLM-R improves performance the amount of benefit it provides is not uniform and is insufficient to prove viable in the very low resource cases. Banyan is able to learn competitive representations consistently across languages, unsupervised and with very little data, providing a viable alternative for cheaply and efficient embeddings for low resource languages.
5.3 Efficiency
Figure 4: Total number of nodes in entangled vs. sentential trees as batch size grows.
Other than its embedding matrix, Banyan only has composition and decomposition functions. Diagonalising these makes them easier to compute and more lightweigh than standard weight matrices, ($2U$ rather than $2U\times U$), achieving a further order of magnitude reduction in parameters compared with the already minimal Self-StrAE.
Secondly, by exploiting entangled tree structure the number of nodes grows at a significantly reduced rate with batch size compared with standard sentential trees (see Figure 4). This is because the number of reused constituent nodes also grows as batch size increases, and entangled trees capture the set of all constituents, which consequently does not grow as drastically. In practical terms, because entangled trees requires fewer nodes, and each node requires two distinct embeddings ($\bar{e}$ and $\underline{e}$) to be held for it, reducing the number of nodes required leads to radical improvements in memory efficiency. Put together, these changes mean that we can train Banyan very quickly as we can use large batches and its small number of parameters ensure quick convergence. On a single Nvidia A40 GPU with a batch size of 1024, Banyan pretrains from scratch in under 50 minutes, meaning that the total cost of pretraining a Banyan model sits at around 30 cents [^2] Cloud computing costs from: https://www.runpod.io/pricing. Free-tier Google Colab users can achieve similar results in about two hours with a smaller batch size. Inference can also be performed on CPU on typical laptops, because the model is so small. Combined with its data efficiency, we believe this provides a promising alternative for low resource languages and communities.
Table 5: Ablations of modelling changes made for Banyan. Higher is better. Results averaged across four random initialisations. Bolded results indicate no standard deviation overlap. Spearman’s $\rho$ is * 100 following convention.
Model STS-12 STS-13 STS-14 STS-15 STS-16 STS-B SICK-R SemRel Score
Standard Trees 31.98 $\pm$ 0.58 53.88 $\pm$ 0.68 37.73 $\pm$ 0.70 55.23 $\pm$ 0.58 $5$$\pm$ 0.47 39.53 $\pm$ 1.61 51.78 $\pm$ 0.29 50.05 $\pm$ 0.92 46.59 $\pm$ 0.43
+ diag functions 35.13 $\pm$ 0.33 56.05 $\pm$ 0.24 40.58 $\pm$ 0.05 58.83 $\pm$ 0.10 56.78 $\pm$ 0.21 44.10 $\pm$ 0.14 53.35 $\pm$ 0.17 52.65 $\pm$ 0.17 49.68 $\pm$ 0.06
++ CE loss 47.10 $\pm$ 1.04 61.85 $\pm$ 1.44 58.60 $\pm$ 1.34 70.45 $\pm$ 0.57 62.45 $\pm$ 0.70 59.50 $\pm$ 0.53 59.00 $\pm$ 0.26 60.33 $\pm$ 0.26 59.91 $\pm$ 0.54
Entangled Trees 38.98 $\pm$ 0.39 61.75 $\pm$ 0.14 43.65 $\pm$ 0.46 58.21 $\pm$ 0.41 55.29 $\pm$ 0.23 46.15 $\pm$ 0.71 53.93 $\pm$ 0.16 52.53 $\pm$ 0.09 51.31 $\pm$ 0.13
+ diag functions 44.15 $\pm$ 0.002 62.80 $\pm$ 0.002 48.30 $\pm$ 0.001 64.60 $\pm$ 0.002 60.30 $\pm$ 0.001 49.80 $\pm$ 0.002 55.14 $\pm$ 0.001 57.70 $\pm$ 0.001 55.23 $\pm$ 0.001
++ CE loss 51.20 $\pm$ 0.007 69.10 $\pm$ 0.002 63.20 $\pm$ 0.004 73.20 $\pm$ 0.002 66.60 $\pm$ 0.002 61.50 $\pm$ 0.002 55.50 $\pm$ 0.003 61.60 $\pm$ 0.002 62.70 $\pm$ 0.001
Table 6: Number of non-embedding parameters.
Model Banyan Self-StrAE RoBERTa (M) All-MiniLM-L12-V2 XLM-R Llama 3.1 Mistral Nemo
Params 14 1072 $\approx$10M $\approx$21M $\approx$85M $\approx$8B $\approx$12B
5.4 Ablations
We have shown that Banyan is more effective than its Self-StrAE predecessor, but to tease apart the impact of proposed changes, we perform ablations on our results from our first set of experiments on English (see Table 5).
The simplest impacts stem from the introduction of the diagonalised composition and decomposition functions. These are sigmoided scalar values with which we multiply embeddings and act similarly to the fast weights of (Ba et al., 2016), decaying in the influence of embeddings further down in the tree on the root representation. This means that the embeddings produced by the model are restricted to conform to the compression order dictated by the structure, and we know from (Opper et al., 2023), that the more we can enforce this constraint the better our representations will end up. Secondly, such simple message passing functions bias the representation space towards informative separability. There has to be some signal from which to perform reconstruction, and all the pressure is now on the representations.
Switching the objective to cross entropy over the vocabulary rather than the contrastive objective used by (Opper et al., 2023) also yields significant benefits. This is likely because the contrastive loss is supposed to be beneficial because it enforces a pressure for representations to be uniformly distributed in space (Wang & Isola, 2020). However, our other modelling changes already push towards this quality. While it is a shame because the contrastive loss is conceptually elegant. It is known to have problems and eventually lead to shortcut solutions (Robinson et al., 2021). Therefore having a more robust objective grounded in data, like the cross entropy over the vocabulary, is actually quite nice.
Finally, changing to entangled trees is also beneficial. The effect is more pronounced before switching to the cross entropy objective, as it removes the issue of false negatives as discussed in Section 4. However, it also is beneficial beyond this. Entangling explicitly allows the model to take advantage of shared constituency structure between complex sequences, as it combines the information from all incoming parent messages. The fact that performance improves using the cross entropy objective shows that explicitly allowing the model to take advantage of such systematicity is useful.
6 Conclusion, Limitations and Future Work
We introduce Banyan, a Self-Structuring AutoEncoder. Banyan’s focus on global, entangled structure and simplified message passing exploits the benefits of structured compositions inherent in language data. It is more effective and efficient than prior work from which we draw three central conclusions.
Firstly, explicitly modelling structured compositions is an effective inductive bias. Table 6 shows the parameters for the structured models versus the baselines. The structured models are far smaller, with tens or thousands of parameters instead of millions or billions. And nonetheless, Banyan is still competitive across several metrics, indicating we have found an efficient learning procedure.
Secondly, we have not yet fully exploited the potential of the inductive bias. Banyan still relies on greedy agglomerative clustering to induce structure. This is effective, but sub-optimal. Future work could learn the structure induction procedure. The type of structure models are exposed to impacts the quality of learnt semantic representations (Opper et al., 2023). So if how we induce structure improves, the model should learn better representations.
Finally, good and cheap embedding models are useful for many applications. For example, the digital humanities need to organise corpora of ancient languages, making it easier for researchers to access texts they need. But these corpora are small, and these languages are unlikely to be present in pretraining corpora of larger models. Banyan provides an efficient solution for producing representations for both these use cases and low resource languages and under represented communities more generally. To conclude, Banyan addresses the problem of efficient learning in low-resource settings.
Impact Statement
This paper represents an effort to enable resource efficient embedding techniques. We hope that our model makes a positive step towards making AI research and technologies more fair, equitable and accessible.
Appendix A The k and u balance
The change to diagonal composition functions allows us to reduce the number of total parameters while maintaining performance. This is because the number of parameters is directly proportional to channel size u. We show ablations for this finding in Table 7. Our findings are similar to those of (Opper & Narayanaswamy, 2024) the smaller the channel size the better the model performs, although in our case we keep things stable between seeds whereas for them when they simplified they faced issues with extreme instability during training. This is thanks to the new message passing functions.
Table 7: Performance Depending on k and u values using new functions. Scores are the average of four random seeds.
k u Lex Score STS Score
4 64 42.9 $\pm$ 0.01 43.5 $\pm$ 0.04
8 32 43.2 $\pm$ 0.02 48.6 $\pm$ 0.01
16 16 47.02 $\pm$ 0.03 62.2 $\pm$ 0.01
32 8 49.2 $\pm$ 0.01 62.9 $\pm$ 0.01
64 4 48.7 $\pm$ 0.01 62.9 $\pm$ 0.01
128 2 49.61 $\pm$ 0.02 62.7 $\pm$ 0.001
256 1 48.7 $\pm$ 0.01 62.9 $\pm$ 0.001